Solve the differential equation dy/dx=(x-y)/(x+y),with y(0)=1 numerically for x=0.1 by Eular's method(five step)

1
Expert's answer
2020-09-03T18:02:22-0400
"Solution"

Let's solve the differential equation by Eular's method. We had the initial value problem:


"\frac{dy}{dx}=\frac{x-y}{x+y}\\
y(0)=1"

Step 1

We'll start at the point "(x_0,y_0)=(0,1)" and use step size of "h=0.01" and proceed for 10 steps. That is, we'll approximate the solution from "t=0\ to\ t=0.1" for our differential equation. We'll finish with a set of points that represent the solution, numerically.


We already know the first value, when "x_0=0" which is "y_0=1" (initial value)


We now calculate the value of the derivative at this initial point. (This tells us the direction to move.)


"\frac{dy}{dx}=f(0,1)=\frac{-1}{1}=-1"


This means the slope of the line from t=0 to t=0.1 is approximately -1


Step 2

Now, for the second step, ( Since "h=0.01", the next point is "x+h=0+0.01=0.01" ), we substitute what we know into Euler's Method formula, and we have:


"y(x+h) \approx y(x)+hf(x,y)\\
y_1 \approx y_0+hf(x_0,y_0)\\
y_1=y(0.01) \approx 1+0.01(-1)=0.99\\"

This means the approximate value of the solution when "x=0.01\ is\ 0.99"

We'll need the new slope at this point, so we'll know where to head next.


"\frac{dy}{dx}=f(0.01,0.99)=\frac{-0.98}{1}=-0.98"


This means the slope of the approximation line from "x=0.01\ to\ x=0.02\ is\ -0.98". So it's a little bit steeper than the first slope we found.


Step 3

Now we are trying to find the solution value when "x=0.02". We substitute our known values:

"y(x+h) \approx y(x)+hf(x,y)\\
y_2 \approx y_1+hf(x_1,y_1)\\
y_2=y(0.02) \approx 0.99+0.01(-0.98)=0.9802\\"

We'll need the new slope at this point, so we'll know where to head next.


"\frac{dy}{dx}=f(0.02,0.9802)=\frac{-0.9602}{1.0002}=-0.96"


This means the slope of the approximation line from "x=0.02\ to\ x=0.03\ is\ -0.96". So it's a little bit steep than the first 2 slopes we found.


Step 4

Now we are trying to find the solution value when "x=0.03". We substitute our known values:


"y(x+h) \approx y(x)+hf(x,y)\\
y_3 \approx y_2+hf(x_2,y_2)\\
y_3=y(0.03) \approx 0.9802+0.01(-0.96)=0.9706\\"

We'll need the new slope at this point, so we'll know where to head next.


"\frac{dy}{dx}=f(0.03,0.9706)=\frac{-0.9406}{1.0006}=-0.94"


This means the slope of the approximation line from "x=0.03\ to\ x=0.04\ is\ -0.94". So it's a little bit steep than the first 3 slopes we found.


Step 5

We continue to calculate the subsequent steps to "x=0.1"

We present all the values up to "x=0.1" in the following table.





(There's no final "\frac{dy}{dx}" ​ value because we don't need it. We've found all the required "y" values.)


We have obtained that the approximation at "x=0.1, \ y=0.909"

The differential equation "\frac{dy}{dx}=\frac{x-y}{x+y}" at "x=0.1" is "0.909--->Answer"